### v2

# Hic Sunt Dracones: Uncovering Dynamical Perturbers Within the Habitable Zone

Abstract
The continuing exploration of neighboring planetary systems is providing deeper insights into the relative prevalence of various system architectures, particularly with respect to the solar system. However, a full assessment of the dynamical feasibility of possible terrestrial planets within the Habitable Zones (HZ) of nearby stars requires detailed knowledge of the masses and orbital solutions of any known planets within these systems. Moreover, the presence of as yet undetected planets in or near the HZ will be crucial for providing a robust target list for future direct imaging surveys. In this work, we quantify the distribution of uncertainties on planetary masses and semi-major axes for 1062 confirmed planets, finding median uncertainties of 11.1% and 2.2%, respectively. We show the dependence of these uncertainties on stellar mass and orbital period, and discuss the effects of these uncertainties on dynamical analyses and the locations of mean motion resonance. We also calculate the expected radial velocity (RV) semi-amplitude for a Neptune-mass planet in the middle of the HZ for each of the proposed Habitable Worlds Observatory target stars. We find that for more than half of these stars, the RV semi-amplitude is less than 1.5 m s-1  rendering them unlikely to be detected in archival RV data sets and highlighting the need for further observations to understand the dynamical viability of the HZ for these systems. We provide specific recommendations regarding stellar characterization and RV survey strategies that work toward the detection of presently unseen perturbers within the HZ.
astrobiology – planetary systems – techniques: radial velocities – planets and satellites: dynamical evolution and stability
1 Introduction
The number of known exoplanetary systems now number in the thousands, enabling statistical studies of exoplanet demographics as well as detail characterization of individual systems (Ford, 2014; Winn & Fabrycky, 2015; He et al., 2019). Improvements in radial velocity (RV) precision combined with long-term survey strategies are gradually filling out the wider-separation system architectures for the nearest and brightest stars, revealing giant planets beyond the snow line (Fischer et al., 2016; Wittenmyer et al., 2020; Fulton et al., 2021; Rosenthal et al., 2021) and providing comparisons to the orbital configuration of the solar system (Horner et al., 2020; Kane et al., 2021). However, there remain significant limitations on how well existing observations can probe for low-mass planets in the Habitable Zone (HZ) of these stars, many of which will serve as targets for future observations that focus on potentially habitable worlds (Kasting et al., 1993; Kane & Gelino, 2012a; Kopparapu et al., 2013, 2014; Chandler et al., 2016; Kane et al., 2016; Hill et al., 2018, 2023). Furthermore, improving the orbital parameter precision of known exoplanets and their host stars is of particular importance for enabling follow-up observations and system analyses (Ford, 2005; Kane et al., 2009). For example, reduced uncertainties in planetary masses are crucial for robust interpretation of atmospheric spectra (Batalha et al., 2017, 2019; Changeat et al., 2020; Di Maio et al., 2023) and interior models (Unterborn et al., 2016; Müller et al., 2020), whilst stellar uncertainties can impact the correct determination of HZ boundary locations (Kane, 2014, 2018) and of a rocky planet’s core mass fraction (Schulze et al., 2021). The continued observations of nearby stars to concisely catalog their planetary inventories is thus a necessary endeavor for enabling subsequent characterization studies of these systems.
A coming evolution of exoplanet detection and exploration is that of directly imaging facilities that will prioritize the detection and characterization of the aforementioned HZ terrestrial planets (Brown, 2015; Barclay et al., 2017; Kane et al., 2018; Kopparapu et al., 2018; Stark et al., 2020; Li et al., 2021). Of particular note are the Habitable Worlds Observatory (HWO) and the Large Interferometer For Exoplanets (LIFE) mission, currently under various development phases for consideration as a means to realize this HZ direct imaging capability (Quanz et al., 2022; Vaughan et al., 2023; Stark et al., 2024). A major task for the development of these missions is the construction and refinement of potential target star lists. An initial list of 164 HWO target stars has been proposed by Mamajek & Stapelfeldt (2024) for discussion and assessment by the community. Science yield calculations based on these proposed HWO targets have been calculated by Morgan et al. (2023) and Tuchow et al. (2024), and the stellar properties have been cataloged in detail by Harada et al. (2024a).
A further consideration is the presence of planets in these systems that may dynamically impede the presence of stable orbits within the HZ. A recent study by Kane et al. (2024) found that 11 of the 30 HWO stars that are known to harbor planets have architectures that eliminate or drastically diminish the ability of those stars to host terrestrial HZ planets. In order to further such analyses for these (and other) stars, continued RV and astrometric measurements must be carried out to both reduce the uncertainties of the derived planetary parameters and detect previously unknown planets. This is especially important for Neptune or sub-Neptune planets within the HZ that, although lying beneath the sensitivity of most archival RV datasets, can still act as gravitational perturbers whose presence has dire consequences for dynamical stability of terrestrial HZ planets.
In this paper, we present the results of an analysis of current planetary parameter uncertainties, their effect on orbital dynamics, and the calculated RV signature of potential dynamical perturbers that may yet be lurking within the HZ of the proposed HWO target stars. To borrow from an old Latin phrase, “Hic sunt dracones”, meaning “Here be dragons”, is an apt phrase when exploring the HZ for objects whose presence may render the region unsuitable for habitable environments. In Section 2, we provide an assessment of the derived uncertainties for planet mass and semi-major axis, both of which are key parameters for dynamical simulations, and their dependence on stellar mass and orbital period measurements. In Section 3, we describe a proposed HWO target list and compare the expected RV semi-amplitudes for a Neptune-mass planet in the center of each star’s HZ, along with the challenges in achieving the necessary RV precision to detect such signatures. Section 4 provides an additional discussion regarding the implications of our results and recommendations for a continued dynamical assessment of stars that are considered as future direct imaging targets. A summary of our work and concluding remarks are delivered in Section 5.
2 Derived Planetary System Uncertainties
To investigate the uncertainties in planetary system properties, we extracted data from the Confirmed Planets Table111doi:10.5281/zenodo.11020 of the NASA Exoplanet Archive (Akeson et al., 2013) as accessed on 2024 August 22 (NASA Exoplanet Archive, 2024). We selected the default parameter set for each planet included in the table, and selected only those planets with valid (non-null) values and uncertainties for the semi-major axis ($a$), planet mass ($M_{p}$ or $M_{p}\sin i$), orbital eccentricity ($e$), orbital period ($P$), and the mass of the host star ($M_{\star}$). This provided 1062 planets that passed our criteria for inclusion in our sample. We note that a small number of the planets in this list have masses derived from transit timing variations ($<4$%) rather than radial velocity data but we restrict the overall considerations of this paper to RV-focused efforts as the transit probability for planets approaching the HZ of sun-like stars falls off quickly (Kane & von Braun, 2008).
2.1 Correlations of Uncertainties
Figure 1: Percentage uncertainties for the planetary masses (left column) and semi-major axes (right column) of our confirmed planet list as a function of their respective parameters (top row; panels A and B), stellar mass uncertainties (middle row; panels C and D), and orbital period uncertainties (bottom row; panels E and F). Planets detected only in RV surveys are denoted as cyan circles, while those detected in both RV and transit surveys are denoted as maroon diamonds. The power law relations between stellar and planetary masses and the semi-major axis, orbital period, and stellar mass (see Equations 2 and 4) are evident as diagonal crowdings of points that set the lower limit in the planet mass and semi-major axis uncertainties. For stars with mass uncertainties above $\sim$20%, the stellar mass uncertainty dominates the planet mass determination, rather than the precision of the RV semi-amplitude measurement. Orbital period uncertainties, however, play only a minor role in the planet mass uncertainties. When calculating the semi-major axis, orbital period uncertainties beyond $\sim$10become the dominant source of error but the stellar mass uncertainty does not produce a distinct boundary identifying when it begins dominating the semi-major axis determination.
From our extracted sample of exoplanets, we focused on the two main planetary properties used as input into dynamical models: the planet mass and semi-major axis. These properties are generally derived from other system observables, and their associated uncertainties are propagated accordingly. In the case of RV observations, the measured semi-amplitude of the RV variations, $K$, is given by
$K=\left(\frac{2\pi G}{P}\right)^{1/3}\frac{M_{p}\sin i}{(M_{p}+M_{\star})^{2/3}}\frac{1}{\sqrt{1-e^{2}}}$ (1)
where the stellar mass, $M_{\star}$, may be determined via relationships derived from binary stars (Torres et al., 2010) or through a combination of spectroscopy (Yee et al., 2017), asteroseisomology (Huber et al., 2017), and spectral energy distribution analyses (Stassun et al., 2018). Assuming $M_{p}\ll M_{\star}$, the planet mass is then calculated from the following relation
$M_{p}=\left(\frac{P}{2\pi G}\right)^{1/3}\frac{M_{\star}^{2/3}K\sqrt{1-e^{2}}}{\sin i}$ (2)
where the orbital period, $P$, and RV semi-amplitude, $K$, are the primary observables. Similarly, Kepler’s third law
$P^{2}=\frac{4\pi^{2}}{G(M_{p}+M_{\star})}a^{3}$ (3)
may be rearranged, assuming $M_{p}\ll M_{\star}$, as follows
$a=\left(\frac{GM_{\star}}{4\pi^{2}}P^{2}\right)^{1/3}$ (4)
to provide an expression for the semi-major axis, $a$.
In Figure 1, we provide plots of the percentage uncertainties for the planetary mass (left column) and semi-major axis (right column) for all of the planets in our sample. Planets detected only in RV surveys are denoted as cyan circles, while those detected in both RV and transit surveys are denoted as maroon diamonds. We adopt an uncertainty definition of the mean of the absolute values of the upper and lower uncertainties on a parameter divided by the value of the parameter itself and then multiplied by 100. The top row plots these uncertainties as a function of the parameters themselves. To provide insight into the dependency of these uncertainties on the other variables in Equation 2 and Equation 4, we also plot the uncertainties as a function of the stellar mass uncertainties (middle row) and orbital period uncertainties (bottom row).
The planet mass distribution shown in panel A exhibits a well-known gap within the range 0.1–0.3 $M_{J}$, first predicted by Ida & Lin (2004) and still a point of discussion as to its observational significance (Mazeh et al., 2016; Bennett et al., 2021; Emsenhuber et al., 2021; Hill et al., 2023). The planet mass uncertainties have a remarkably wide range of 0.2–622%, but it should be noted that the planet mass uncertainty outlier at 622% is the circumbinary planet Kepler-47 b where the mass uncertainties are dominated by a relatively unconstrained upper mass limit (Orosz et al., 2019). The planet mass uncertainty distribution has mean and median values of 16.0% and 11.1%, respectively. Panel C shows that, for stellar mass uncertainties beyond $\sim$20%, the stellar mass uncertainty dominates the planet mass determination, below which it is the precision of the RV semi-amplitude measurement that dominates. The orbital period uncertainties shown in panel E demonstrate that the precision of the orbital period measurements play only a minor role in the planet mass uncertainties.
Figure 2: Percentage uncertainties for the planetary masses (left column) and semi-major axes (right column) of our confirmed planet list as a function of host star brightness. Planets detected only in RV surveys are denoted as cyan circles, while those detected in both RV and transit surveys are denoted as maroon diamonds. The impact of transit surveys, which can more efficiently detect planets around fainter stars, is notable but neither planetary uncertainty shows a significant trend with host star magnitude.
The semi-major axes and associated uncertainties are shown in panel B of Figure 1. As for the planet mass uncertainties, the semi-major axis uncertainties span several orders of magnitude, falling in the range 0.005–60%, with mean and median values of 4.2% and 2.2%, respectively. For the planets detected in RV surveys only (cyan circles), the data in panels B, E, and F, show a clear capability of the RV surveys to detect longer period planets relative to the transit detections as expected given the decrease in transit probability with increasing orbital period (Kane & von Braun, 2008). The semi-major axis uncertainties show a distinct rise at separations higher than $\sim$5 AU, where the uncertainties become dominated by the those uncertainties associated with the orbital period rather than stellar mass. This correlation is validated by the data distributions shown in panels D and F, where indeed it can be seen that the uncertainties on semi-major axis correlate broadly with those of stellar mass, but correlates strongly with those of orbital period in the high uncertainty regime. The panels of Figure 1 thus provide a broad overview of the relative contributions toward the planet mass and semi-major axis uncertainty budgets, and highlight the importance of improving characterization of host stars and their mass uncertainties. We further provide the range of percentage uncertainty values for each parameter discussed here, along with the mean, median, standard deviation, median absolute deviation (MAD), and interquartile range (IQR) values, in Table 1. The values shown in the the table are for all planet masses, and also for the planet mass groups on either side of 0.1 $M_{J}$ to emphasize the effects of planet mass on the measurement precision. Note that the high uncertainty for semi-major axis and orbital period for the high mass group ($>0.1$ $M_{J}$) is the result of the observational bias for high mass planets at large separations.
Table 1: Parameter percentage uncertainty statistics.
All planet masses
Parameter Range Mean Median Std. Dev. MAD IQR
% Unc. in $M_{p}$ 0.2–622.46 16.0 11.1 24.27 5.29 11.8
% Unc. in $M_{\star}$ 0.76–181.57 8.4 5.7 10.22 2.93 8.1
% Unc. in $P^{\dagger}$ 0.0–90.16 2.8 0.1 9.35 0.1 0.8
% Unc. in $a$ 0.005–60.0 4.2 2.2 6.16 1.36 3.0
$M_{p}\leq 0.1$ $M_{J}$
Parameter Range Mean Median Std. Dev. MAD IQR
% Unc. in $M_{p}$ 1.9-622.46 21.9 15.0 38.4 5.8 12.7
% Unc. in $M_{\star}$ 0.76-23.33 5.5 4.5 3.63 1.77 3.6
% Unc. in $P^{\dagger}$ 0.0-6.62 0.2 0.009 0.63 0.01 0.1
% Unc. in $a$ 0.005-47.12 2.7 1.7 3.51 0.84 2.4
$M_{p}>0.1$ $M_{J}$
Parameter Range Mean Median Std. Dev. MAD IQR
% Unc. in $M_{p}$ 0.2–195.45 13.7 9.5 15.0 4.7 10.5
% Unc. in $M_{\star}$ 0.76–0.76 9.6 6.6 11.64 3.53 8.2
% Unc. in $P^{\dagger}$ 0.0–0.0 3.9 0.3 10.84 0.31 1.4
% Unc. in $a$ 0.01–0.01 4.8 2.6 6.84 1.5 3.3
footnotetext: $P$ uncertainties of zero correspond to uncertainties $<10^{-4}$ days.
A further consideration is the effect of host star brightness on measurement signal-to-noise and subsequently the uncertainties. Shown in Figure 2 are the percentage uncertainties for the planetary masses (left column) and semi-major axes (right column) of our exoplanet sample as a function of host star brightness. As for Figure 1, the data are separated by those detected via RV surveys only (cyan circles) and those detected via RV and transit surveys (maroon diamonds). Although transit surveys can more efficiently detect planets around fainter stars than RV surveys, it is notable that the distribution of uncertainties for planet mass and semi-major axis do not substantially differ as a function of host star brightness. This may be due to the fact that transiting planets provide tight constraints on the planet’s period and transit epoch thereby reducing the number of free parameters in the fit and allowing for more precise mass and semi-major axis uncertainties. Transiting planets also allow for more intentional RV coverage across the planet’s orbital phase which results in more accurate and precise masses (Burt et al., 2018; Lam et al., 2024).
2.2 Effects on Dynamical Simulations
As described in Section 2.1, an impetus for improving the planet mass and semi-major axis values lies in their critical role in assessing the dynamical stability and evolution of planetary systems (Maire et al., 2023). Orbital dynamics has proven to be a useful tool in predicting the possible locations of additional planets in known systems (Barnes & Raymond, 2004; Rivera & Haghighipour, 2007; Kopparapu et al., 2009; Agnew et al., 2019), as well as assessing their dynamical packing (Fang & Margot, 2013; Obertas et al., 2017; Kane et al., 2020; Kane, 2023a; Obertas et al., 2023). Dynamical simulations have also been applied to evaluating dynamical stability in the HZ, with application to terrestrial planet climates and the sustainability of temperate conditions (Williams & Pollard, 2002; Dressing et al., 2010; Kane & Gelino, 2012b; Carrera et al., 2016; Kane & Blunt, 2019; Kane et al., 2022; Li et al., 2022; Vervoort et al., 2022). Such work relies upon the robustness of reported planetary system parameters, both in terms of the measurement precision and accounting for other possible planets and interactions (Beaugé et al., 2008; Wittenmyer et al., 2019).
The precise effects of system uncertainties on dynamical analyses are highly dependent upon the system architecture under consideration (Hadden & Lithwick, 2018; Kane, 2023b). By way of example, consider the HD 141399 system, which contains four giant planets with a diversity of star–planet separations (Vogt et al., 2014). The dynamical analysis of the HD 141399 system provided by Kane (2023b) revealed the complex mean motion resonance (MMR) structure of the system, and the consequences for dynamical stability within the HZ. Many of the adjacent primary MMR locations are separated by a distance of only 10%, such as the 3:1 (0.201 AU) and 5:2 (0.227 AU) MMR locations for planet b. As shown in Section 2.1, while the median semi-major axis uncertainty is only 2.2%, values can be as high as 60%. These semi-major axis uncertainties can have significant consequences for the correct diagnosis of dynamical aspects, such as the apsidal behavior of orbital evolution (libration or circulation) near MMR locations Barnes & Greenberg (2006); Kane & Raymond (2014) and can ultimately lead to incorrect conclusions regarding the long-term stability prognosis of planetary systems.
3 Perturbers Within the Habitable Zone
In this section, we discuss the importance of detecting additional gravitational perturbers in the HZ, and the RV precision required to achieve this for HWO targets.
3.1 Habitable Worlds Observatory Targets
Figure 3: The RV semi-amplitude of a Neptune-mass planet in the middle of the OHZ for each of the stars on the Mamajek & Stapelfeldt (2024) HWO target list, shown as a function of stellar effective temperature (left panel) and as a histogram (right panel). The vertical dashed line indicates the Kraft break at Teff = 6250 K, beyond which precision RV measurements become increasingly challenging due to the rapid rotation of the stars and the associated broadening of the stars’ absorption lines. The majority these hypothetical HZ Neptunes would have RV semi-amplitudes below 1.5m s-1 making them challenging detections, especially at the 100+ day orbital periods of these HZs. Long term, dedicated EPRV survey efforts would be required to detect these potential perturbers and determine what portions of each star’s HZ remains dynamically viable for Earth analog planets. For the hottest stars, similar surveys via precision astrometry could provide complementary insights into the star’s HZ inhabitants.
The recent work by Kane et al. (2024) demonstrated the importance of understanding the full dynamical effects of system architectures on the possible presence of planets in the HZ. Here, we use the proposed HWO target list provided by Mamajek & Stapelfeldt (2024) to calculate the signature of an additional, Neptune-mass, planet within the HZ (described in Section 3.2) for all 164 target stars. The stars are all relatively bright, with a V magnitude range of 0.0–7.4, and span a broad range of spectral types, with an effective temperature range of 3601–6990 K.
For each star, we calculated the HZ boundaries using the stellar parameters provided by Mamajek & Stapelfeldt (2024) and the prescription described by Kopparapu et al. (2013, 2014). This methodology adopts the products from 1D Earth-based climate models to determine the radiative balance for which surface liquid water is retained for a planet with an Earth-based atmospheric composition. The HZ is considered to contain two primary regions: the conservative HZ (CHZ) and the optimistic HZ (OHZ). The inner edge of the CHZ is the boundary where water vapor saturates the stratosphere and the oceans evaporate entirely, and the outer edge of the CHZ is the boundary where the CO2 greenhouse effect maximizes due to increased Rayleigh scattering (Kopparapu et al., 2013; Kane et al., 2016). The boundaries of the OHZ are empirically determined boundaries that extend those of the CHZ, based on evidence regarding the possible retention of surface water early in the histories of both Venus and Mars (Baker, 2001; Kane et al., 2014; Way et al., 2016; Orosei et al., 2018; Kane et al., 2019a; Kane & Byrne, 2024). These calculated HZ boundaries were used in our subsequent analysis for placing additional planets around the HWO stars.
3.2 Required Radial Velocity Observations
There are a variety of planet mass and orbits that can cause significant impact orbital stability within the HZ, from those whose orbits lie entirely within the HZ to giant planets on much longer period but more eccentric orbits (e.g., Kane & Blunt, 2019). Here, we consider the simple case of a Neptune-mass planet ($M_{p}=17$ $M_{\oplus}$) on a circular orbit in the middle of the HZ. This particular configuration was chosen based on the recent re-analysis of the GJ 357 system by Kane & Fetherolf (2023), which showed that even a 10 $M_{\oplus}$ mass planet in the HZ can render a substantial fraction of that region unstable for other terrestrial planets. For each of the 164 proposed HWO targets stars (see Section 3.1), we calculated the OHZ boundaries and the RV semi-amplitude for a Neptune-mass planet that lies halfway between them. The results of these calculations are shown in Figure 3, both as a function of the host star effective temperature (left panel) and as a histogram (right panel). As expected, the RV semi-amplitude decreases with increasing stellar effective temperature, since both the stellar mass and semi-major axis of the HZ boundaries increase in that direction. There are two M dwarfs (GJ 411 and GJ 887) for which the RV semi-amplitude exceeds 6 m s-1, but the vast majority have a RV semi-amplitude that is less than 5 m s-1.
[…]
Recovering long period, small RV semi-amplitude signatures remains challenging in the face of stellar activity and correlated noise (Narayan et al., 2005; Kane et al., 2009; Wittenmyer et al., 2013; Fischer et al., 2016; Langellier et al., 2021; Luhn et al., 2023). For example, the community-wide RV challenge conducted by Dumusque et al. (2017) used data acquired with the High Accuracy Radial velocity Planet Searcher (HARPS) spectrograph (Pepe et al., 2000), into which planetary signatures were injected. The results of that challenge indicate that, for a star with 1.5 m s-1stellar jitter, a planetary signature with a RV amplitude of 1.5 m s-1 would require $\sim$60 HARPS observations to achieve a robust detection, depending on the brightness of the host star, the observing cadence adopted, and the precision of the specific instrument used. Indeed, recent studies of archival RV measurements by Laliotis et al. (2023) and Harada et al. (2024b) demonstrated that the RV sensitivity is not yet sufficient to extract the signature of Saturn-mass planets that may be present in the HZ for many of the proposed HWO target stars. Additional effort will be required for a more detailed assessment of chromospheric activity (Cale et al., 2021; Isaacson et al., 2024), particularly for the coolest host stars where the HZ orbits and overlap with stellar rotation periods (Vanderburg et al., 2016). Although a full diagnosis of the specific number of observations and optimum scheduling, reduction, and analysis strategies required for all HWO stars is a complex and multi-dimensional problem, it is clear that significant observational resources are required to adequately uncover the signatures of potential gravitational perturbers that yet remain in or near the HZ of those systems. We provide specific recommendations toward achieving the goal of detecting such planets in Section 5.
4 Discussion
There are several additional points of consideration to add to the details of this paper. Since the discussion here has been largely focused on the planetary signatures and uncertainties that arise from RV observations, there remains the problem of an unknown orbital inclination in the majority of these cases. The RV semi-amplitude calculations of Section 3.2 assume an edge-on inclination, which will not be true in the vast majority of cases, resulting in more stringent precision requirements to detect the correspondingly smaller amplitude planetary RV signatures. Orbital dynamics for multi-planet systems may be used to place lower limits on the inclination of the planetary orbits, and subsequent effects on the HZ (Kane & Gelino, 2014; Korolik et al., 2023). The coplanarity of the orbits within a system is difficult to determine for non-transiting systems (Fang & Margot, 2012; Ballard & Johnson, 2016), however, and will also influence the calculated dynamics of the HZ (Kane, 2016). A further pathway toward resolving the inclination ambiguity is the use of astrometric observations, in particular those of the Gaia mission (Perryman et al., 2014; Gaia Collaboration et al., 2021; Feng et al., 2021) in combination with RV data (Brandt et al., 2019; Kiefer et al., 2021; Winn, 2022). Such methodology may only be applied for those stars that are fainter than the Gaia bright limit of $G\approx 3$ (Lindegren et al., 2018). Though Gaia is unlikely to reveal planets within the HZ, the constraints placed on the inclination of outer, giant planetary companions can provide valuable insight into the inclination and dynamics of inner planets within the system. The use of Gaia data has the additional benefit of greatly improving stellar mass uncertainties (Stassun et al., 2017; Fulton et al., 2017), and thus increasing the reliability of dynamical simulations, as discussed in Section 2. However, the inevitable outcome of such observations will be to reveal that the true planetary masses are indeed larger than the minimum masses inferred from the RV data alone, further inhibiting the dynamical real estate available with the HZ of target systems.
The calculations presented here assume a circular orbit for the Neptune-mass planet in the HZ but there are a broad range of orbital architectures that may influence this region. For the known systems examined by Kane et al. (2024), the majority of HZ instability cases were due to giant planets in eccentric orbits that pass in or near the HZ. However, the discussion here is concerned with smaller, as yet undetected, planets that can still render HZ orbits for terrestrial planets unviable. The dynamical analysis of the GJ 357 system by Kane & Fetherolf (2023) demonstrated that even a relatively small eccentricity of 0.1 can have a major effect on the exclusion of other planetary orbits nearby. Thus, the results presented here may be considered optimistic in the context of circular orbits being the dynamically most favorable scenario.
Exoplanet characterization via RV observations will also be a crucial component of ephemeris refinement to optimize the use of direct imaging facilities (Kane, 2013; Kane et al., 2019b; Dulz et al., 2020). These techniques can also be applied to determining the presence of possible planetary and/or stellar companions to target stars at wide separations (Crepp et al., 2012; Wittrock et al., 2016, 2017; Kane et al., 2019b; Dalba et al., 2021; Rickman et al., 2022; Howell et al., 2024). If such high-mass companions are present, they can play a powerful role in the dynamical evolution of the system, and can also contribute toward the scattering of volatiles from beyond the snow line, that will ultimately build the volatile inventory of inner terrestrial planets (Raymond et al., 2004; Ciesla et al., 2015; Marov & Ipatov, 2018; Ogihara et al., 2023; Kane & Wittenmyer, 2024).
Finally, it is worth highlighting the importance of this work in the design considerations for the next generation of direct imaging missions. Revealing system architectures that can effectively impede the presence of HZ planets within a system requires that replacement targets be identified in order to avoid compromising the predicted mission yield (Kopparapu et al., 2018; Stark et al., 2019). Delving into a fainter population of stellar targets may necessitate an increased aperture to achieve the required measurement signal-to-noise. Alternatively, the use of a starshade as an external occulter, though technologically challenging, can dramatically reduce the contrast ratio requirements for a successful observation of a HZ terrestrial planet (Turnbull et al., 2012; Hu et al., 2021; Turnbull et al., 2021).
5 Conclusions
The preparation for high contrast, space-based direct imaging missions requires a significant investment in the stellar characterization of potential targets. Inadequate precursor studies carry the risk of spending valuable mission time observing targets that are unsuitable for achieving the stated goals of the mission. Where the goal is to directly image terrestrial planets within the HZ, the use of ground-based RV facilities should be sufficiently leveraged to ensure the success of the mission. One such critical task is assessing the viability of system HZ regions to dynamically harbor a terrestrial planet.
Our work here has demonstrated the limitations imposed by the current state of uncertainties for fundamental system parameters, including the measured property of orbital period, and the derived properties of stellar mass, planet mass, and semi-major axis. A major source of uncertainty originates from the stellar mass uncertainties, which can depend upon the utilized methodology (Basu et al., 2012; Valcarce et al., 2013; Young et al., 2014). For our sample of 1062 planets, the median planet mass and semi-major axis uncertainties are 11.1% and 2.2%, respectively, posing a challenge for subsequent dynamical work. The main dynamical concern is the possible presence of planets within the HZ whose existence serves to effectively exclude the terrestrial planets that missions, such as LIFE and HWO, aim to observe.
We show that a hypothetical Neptune-mass planet in the middle of the OHZ of the HWO stars induces an RV semi-amplitude less than 1.5 m s-1for the majority of the ExEP HWO target list. Signals of this size remain challenging for modern RV facilities to detect, especially at 100+ day orbital periods of the HZs for most potential HWO target stars, and and will require long-term, dedicated, observational campaigns to uncover. We therefore recommend two main community actions to address the issues raised in this work.
First, we recommend that the exoplanet community invest in an update of the stellar parameters for all known exoplanet host stars using a self-consistent and homogeneous analysis. One potential approach is to fit the stars’ spectral energy distributions, combined with precise distance estimates from Gaia DR3, and then using Bayesian Model Averaging to incorporate information from multiple stellar models (see, e.g., Vines & Jenkins, 2022; Harada et al., 2024c). Alternatively, high resolution spectra for these stars can be used to derive spectroscopic stellar parameters, either via comparisons to a grid of spectral templates (e.g., Petigura et al., 2017) or via direct spectral synthesis that computes radiative transfer through a model stellar atmosphere (e.g., Valenti & Piskunov, 1996; Soto & Jenkins, 2018). Where possible, the spectroscopically derived parameters should be used as constraints to interpolate upon stellar isochrones (e.g., Demarque et al., 2004; Choi et al., 2016) to derive additional mass, radius, and age ranges for the stars. As most exoplanet characteristics are derived parameters whose value and uncertainty rely on our knowledge of the host star’s mass and/or radius, updates to the stellar parameters will then need to be propagated through to the planet parameters. A full re-analysis of the original time series data that led to the detection of the planet, either the transit photometry or RV time series, would provide the most cohesive view of each system. However, an intermediate option would be to take the measured quantity (either the transit depth or the RV semi-amplitude) and re-derive the corresponding exoplanet radius or mass measurement using the updated stellar radius or mass.
Second, we recommend the initiation of dedicated RV surveys that target the list of the likely HWO direct imaging targets and search for evidence of disruptive planets. Recent simulation efforts from Luhn et al. (2023) and observational efforts of the sun from Klein et al. (2024) and Ford et al. (2024) suggest that the detection of HZ Neptunes and even super-Earths at the $\sim$10 Earth mass level is attainable with modern RV instruments. The application of a Fischer information metric that incorporates information about correlated stellar noise (see, e.g., Cloutier et al., 2018; Gupta & Bedell, 2024; Lam et al., 2024) may provide an efficient way to survey the HZs of these stars. The design and execution of such a survey will likely require multiple teams working in coordination, as the HWO targets span the entire sky and cannot be comprehensively observed from a single facility. Such a survey will also require a multi-year investment of observing resources as the orbital periods of these HWO HZs are 100s of days long, and establishing the stability of an RV signal over multiple orbits is often a requirement for confirming its Keplerian nature, particularly for those RV semi-amplitudes near the 1.5 m s-1peak described in Section 3.2. As a first step, we therefore recommend that the exoplanet community determine the most effective and efficient way to conduct such a survey and compare the resulting observational needs with the current suite of accessible RV facilities to determine what additional resources / observing support may be required. Once such a plan has been developed and vetted by the community, execution should begin as soon as possible to ensure that the outer regions of the HZ are thoroughly investigated in time to provide insights to HWO’s architectural design and science operations. The detection of these dynamical perturbers within the HZ will not only guide the target selection of direct imaging missions but may also guide the mission architectures if such perturbers are found to be common place, thus requiring an increase in mission target lists and a corresponding increase in mission aperture or duration. The results of such a survey would also provide fundamental demographics information that informs the occurrence of potentially habitable planets.
Acknowledgements
We acknowledge support from the NASA Astrophysics Decadal Survey Precursor Science (ADSPS) program under grant No. 80NSSC23K1476.
Part of this research was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (NASA).
This research has made use of the NASA Exoplanet Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. This research has also made use of the Habitable Zone Gallery at hzgallery.org. The results reported herein benefited from collaborations and/or information exchange within NASA’s Nexus for Exoplanet System Science (NExSS) research coordination network sponsored by NASA’s Science Mission Directorate.
We follow the https://doi.org/10.5281/zenodo.10161527 (catalog guidelines) of NASA’s Transform to OPen Science (TOPS) mission for our open science practices. The software used to execute the analyses in this paper and to generate the corresponding figures is preserved, alongside the input data sets, on Zenodo at this 10.5281/zenodo.13363032 (catalog DOI).